OpenStudy (anonymous):

standard form: 8/7+i

OpenStudy (saifoo.khan):

\[\frac{8}{7+i}\]Like that?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

i got 56-8i/50 idk what to do after

OpenStudy (anonymous):

(56-8i)/8

OpenStudy (anonymous):

why over 8?!

OpenStudy (anonymous):

\[\frac{8}{7+i}*\frac{7-i}{7-i}=\frac{56-8i}{49+1}=\frac{28}{25}-\frac{4i}{25}\]

OpenStudy (anonymous):

\[\sqrt{2x+3}-\sqrt{x+1}=1\]

OpenStudy (anonymous):

thank you mandolino!!

OpenStudy (anonymous):

you were right... standard form means to separate the real and imaginary parts: a + bi